OpenStudy (anonymous):

how do I find what x equals. (6,-1) with the equation y+(-1)=m(x-6)

OpenStudy (dakid88):

ou have to note that y=mx+b is a linear equation, so the graph is a straight line. In order to build the graph, you have to plug in values for x, and find the values for y. For tracing the line, it's necessary to have two points. That means having 4 coordinates. You know that a point is defined by its coordinate (x,y). y=5x-6 For x = 0 y = 5*0 -6 y = -6 The point has the following coordinates: (0,-6) For x = -1 y = 5*(-1) -6=-5-6=-11 The point has the following coordinates: (-1,-11) Now, in order to graph, trace the system of coordinates, which means two rectangular axes. ox is perpendicular to oy. Choose an unit of measure (for example, if you're working on a math sheet, a unit of measure could be the square from the page) The negative values, on ox axis, are on the left side and the positive values are on the right side. In the middle of the axis is origin, which has coordinates (0,0). The negative values, on oy axis, are above origin, and negative values are below origin. For the point (-1,-11), on ox axis, number 1 square on the left side and, for oy axis, number 11 squares below origin.